# Reconstruct, Inpaint, Finetune: Dynamic Novel-view Synthesis from Monocular Videos

### v1

@@ Abstract
Figure 1: We present CogNVS, a video diffusion model that enables novel-view synthesis of dynamic scenes. Given an in-the-wild monocular video of a dynamic scene, we first reconstruct the scene, render it from the target novel-view and inpaint any unobserved regions. Because CogNVS can be pre-trained via self-supervision, it can also be test-time-finetuned on a given target video, enabling it to zero-shot generalize to novel domains. Our simple pipeline outperforms almost all prior state-of-the-art for dynamic novel-view synthesis. We show outputs from CogNVS from two unseen videos; a generated video above, and a real-world video below.
@@ 1 Introduction
Rapid advances in static 3D scene representations [1, 2] have paved the way for spacetime understanding of the dynamic world. This has enabled photorealistic content creation and immersive virtual reality applications. In this work, we focus on the problem of novel-view synthesis from casually-captured monocular videos of dynamic scenes.
@@
Prior work on dynamic view synthesis addresses this task from two extremes. The first class of methods “test-time” optimize a new 4D representation from scratch for every new test video. While this ensures physically-plausible scene geometry, careful choices in modeling scene motion – in the form of an independent deformation field, or learnable temporal offsets – have to be made [3, 4, 5]. More importantly, it can take on the order of hours to optimize and render a novel-view video. An attractive alternative is to train large feed-forward video models directly for view synthesis [6, 7]. While inference on such models is dramatically faster (on the order of milliseconds), the resulting renderings often are not as accurate as their test-time optimized counterparts. From a pragmatic perspective, such models need to be trained on mega-scale multi-view training data, which is difficult to obtain for dynamic scenes.
@@
addresses the above challenges by decomposing the problem of dynamic view-synthesis into three distinct stages. First, we lean on the success of non-rigid structure from motion [8, 9, 10] approaches that produce reconstructions of visible scene regions, sometimes known as "2.5D" reconstructions (since occluded regions are not reconstructed). We point out that such reconstructions can be trivially produced for casual mobile videos captured with depth sensors and egomotion [11]. When such reconstructions are rendered from a target novel view, previously-hidden regions will not not be rendered. To “inpaint” these regions, we train a 2D video-inpainter – CogNVS – by fine-tuning a video diffusion model (CogVideoX [12]) to condition on the partially-observable novel-view pixels. Importantly, we allow CogNVS to also update the appearance of previously-visible pixels, allowing our pipeline to model view-dependent (dynamic) scene effects.
The key insight of our work is that CogNVS can be trained on any 2D videos via self-supervision. However, rather than training our inpainter with random 2D masks, we make use of 3D multi-view supervision that better captures 3D scene visiblity, similar to prior art [13]. Specifically, given a 2D training video, we first reconstruct it (with an off-the-shelf method such as MegaSAM) and then render the reconstruction from a random camera trajectory. This rendering is used to identify co-visible pixels from the source video that remain visible in the novel views. This original source video and its co-visible-only masked variant can now form a training pair for 3D-consistent video inpainting. Importantly, because such a training pair does not require ground-truth 3D supervision, CogNVS can be trained on diverse in-the-wild 2D videos. We use dynamic scenes from TAO [14], SA-V [15], Youtube-VOS [16], and DAVIS [17]. Equally as important, we use the same paradigm to test-time finetune CogNVS on the test video-of-interest. We show that this allows our pipeline to “zero-shot” generalize to test videos that were never seen during training. We argue that our test-time finetuning of 2D diffusion models can be seen as the “best-of-both-worlds”, by leveraging large-scale training data (for data-driven robustness) and test-time optimization (for accuracy).
@@ 2 Related Work
has seen recent advancements with the rise of implicit scene representations like NeRFs [1] and Gaussian primitives [2, 18, 19]. We have seen widespread efforts in scaling these representations to model larger scenes [20, 21, 22], making them faster to fit [23, 24, 25, 2, 26, 27, 28, 29], anti-aliased [30, 31, 32, 33, 34], and extend to representing dynamic scenes [35, 36, 37, 23]. The most popular paradigms have been the adoption of dynamic NeRFs [1, 37, 38, 39] and deformable Gaussian primitives [2, 5, 40, 41] for modeling scene dynamics, apart from using voxel grids [42, 43] or learnable tokenzation [44]. Most approaches need multi-view posed videos as input, and only recently monocular view synthesis has gained traction [45, 46, 3, 47]. However, each of the aforementioned approaches have to be test-time optimized separately for every new test video, are slow to optimize and yet fail to recover highly-detailed dynamic scene content [46]. Moreover, there is no focus on predicting the unobservable scene content, which is exacerbated by benchmarking metrics that only evaluate co-visible pixels [11] in training and inference views and therefore encourage benchmarking on novel views that are not too far apart from the training views. Our approach instead reformulates dynamic view synthesis as an inpainting task, which specifically focuses on generating parts of the scene that were occluded from the training views, thereby facilitating extreme novel view synthesis for dynamic scenes. Our large-scale pretraining for feed-forward novel-view inpainting enables data-driven robustness.
@@
Figure 2: CogNVS overview. During training (left), given a 2D source video (in blue) of a dynamic scene, we first reconstruct the scene using off-the-shelf monocular reconstruction algorithms like MegaSAM [10] to obtain the 3D scene geometry, $\mathcal{G}_{\rm src}$ and camera odometry, $\mathbf{c}_{\rm src}$. We then sample a set of arbitrary camera trajectories $\{\mathbf{c}_{1},\cdots,\mathbf{c}_{N}\}$ to simulate plausible occluded geometries, $\{\mathcal{G}^{\rm cov}_{{\rm src},1},\cdots,\mathcal{G}^{\rm cov}_{{\rm src},N}\}$ which when rendered from original camera trajectory, $\mathbf{c}_{\rm src}$ produces a mask of source pixels that are co-visible in the sampled trajectory (in orange). The source video and its masked variant produce a self-supervised training pair for learning CogNVS, our video inpainting diffusion model (visualized in Fig. 3). At inference (right), we finetune CogNVS on the given input sequence by similarly constructing self-supervised training pairs. The final novel-view is then generated using the finetuned CogNVS in a feed-forward manner.
@@
approaches have emerged [48, 49, 50, 51, 52, 53, 54] which train for view synthesis in a feed-forward manner with large-scale data. One class of methods is based on transformer architectures, more often than not trained with multi-view supervision and rendering in the loop [7, 6, 55, 56, 57, 58]. Another class of methods reformulate novel-view synthesis as a conditional generation task and use diffusion-based generative architectures [53, 59] for the same. Initially, the focus was on developing data-driven pipelines for static novel-view synthesis [53, 60, 61, 52, 6, 56, 62, 63], or exploiting data-driven priors [64, 65, 66, 67, 68, 69, 70, 71], using multi-view posed image inputs. However, the focus is now shifting to dynamic view synthesis of casually captured videos in an unconstrained setting [48, 50, 51, 54, 72, 73]. These approaches allow a greater level of hallucination of unseen scene components which instills the capability of view synthesis for camera poses that are far apart from the training views. We fit into this setting. While the data-driven learning provides faster inference times and a broad generalization, it compromises on 3D geometric accuracy and physical-plausibility which introduces unrealistic artifacts in the synthesized outputs (e.g., objects suddenly exist or cease to exist) [51]. In this work, we highlight that test-time finetuning is crucial to preserving the 3D geometry of the scene and reducing implausible artifacts.
@@
is a long-standing paradigm to curb distribution shifts in machine learning algorithms and improve their generalization. It’s origin lies in early-age algorithms for optical character recognition [74] and text classification [75], where the algorithm adjusts itself after observing the test data. A decade back, it popularly resurfaced for super-resolution [76] where learning to super-resolve an image was achieved by downsampling and super-resolving the test image. Domain generalization approaches for vision [77, 78, 79, 80, 81, 82] soon took inspiration from this breakthrough and recently, chain-of-thought prompting [83] and general LLM reasoning [84] in natural language processing adapted this paradigm. The most recent adoption was seen in 4D reconstruction and tracking [85], and we similarly explore this paradigm further in our work.
@@ 3 Method
Given a monocular video of a dynamic scene, $\mathbf{V}_{\rm src}=\{\mathbf{V}_{\rm src}^{t}\}_{t=1}^{T}$, we want to generate a novel view of the observed scene, $\mathbf{V}_{\rm nvs}=\{\mathbf{V}_{\rm nvs}^{t}\}_{t=1}^{T}$ from a target camera pose. As discussed (c.f. Fig. 2), we achieve this by decomposing the task into three distinct stages – (1) obtain an off-the-shelf reconstruction of the observed scene over time, (2) render the scene from the novel views and inpaint the non-co-visible regions, and (3) curb the train-test distribution shift with test-time finetuning. We now descibe each of the stages in detail.
@@ 3.1 Dynamic view synthesis as structured inpainting
We use off-the-shelf SLAM frameworks, like MegaSAM [10], to obtain a reconstruction of the given scene. Formally, let the underlying 3D structure of the world as observed by $\mathbf{V}_{\rm src}$ be represented by, $\mathcal{G}_{\rm src}=\{\mathbf{X}_{\rm src}^{t}\}_{t=1}^{T}$, where $\mathbf{X}_{\rm src}^{t}$ are the evolving 3D primitives (points, Gaussians, etc.) across time, $t$. Any physical properties of the primitives are omitted from this discussion for simplicity. Let the recovered camera poses from which $\mathbf{V}_{\rm src}$ was observed be, $\mathbf{c}_{\rm src}=\{\mathbf{c}_{\rm src}^{t}\}_{t=1}^{T}$, where $\mathbf{c}$ denotes a camera pose and is formulated as, $\mathbf{c}=(\mathbf{R},\mathbf{t})\in\mathrm{SE}(3)$ lie group. The source video $\mathbf{V}_{\rm src}$ can be obtained by using a rendering function $\mathcal{R}$ as,
@@
For obtaining $\mathbf{V}_{\rm nvs}$, we note that a subset of 3D primitives that must be visible from $\mathbf{c}_{\rm src}$, are already available in the reconstructed scene geometry, $\mathcal{G}_{\rm src}$. Therefore, a partial observation of the world in the form of co-visible pixels [11] from novel views, $\mathbf{c}_{\rm nvs}=\{c_{\rm nvs}^{t}\}_{t=1}^{T}$, can be rendered as follows,
@@
At this point, the novel view synthesis is incomplete, and all missing regions have to be generated. To this end, we train a conditional video diffusion model, CogNVS (denoted by $\epsilon_{\theta}$) built on top of a recently proposed transformer-based video diffusion model [12]. CogNVS takes in the partially observed novel view video and generates an inpainted novel-view of the scene. The overall CogNVS pipeline first employs a 3D causal VAE to compress the conditioning $\mathbf{V}_{\rm nvs}^{\rm cov}$ and target novel-view $\mathbf{V}_{\rm src}$ into latent representations $\mathbf{z}_{\rm cond}=\mathcal{E}(\mathbf{V}_{\rm nvs}^{\rm cov})$ and $\mathbf{z}_{0}=\mathcal{E}(\mathbf{V}_{\rm src})$ respectively, enabling efficient training while preserving temporal coherence and photometric fidelity. Here, $\mathcal{E}$ is the VAE encoder. Gaussian noise is then added to the target latent $\mathbf{z}_{0}$, and the resulting noisy latent is concatenated with the conditional latent $\mathbf{z}_{\rm cond}$. This joint representation is passed through a self-attention transformer equipped with 3D rotary positional embeddings (3D-RoPE) [86] and adaptive layer normalization, which predicts the added noise. The training objective follows a score matching formulation:
@@ 3.2 Data generation for self-supervised training
Figure 3: Self-supervised training data generation. To curate a large training set for video inpainting, we first reconstruct an input source 2D video (in blue) with an off-the-shelf monocular SLAM system. After reconstruction, we randomly sample $N$ pairs of ‘start’ and ‘end’ camera poses around a spherical region, $\mathcal{S}$ of the estimated camera pose in the given 2D video. $\mathcal{S}$ is bounded by a predefined deviation in the spherical coordinate axes, similar to a prior work [52]. We sample a ${\rm SE(3)}$ camera trajectory that interpolates the start and end poses while looking at the center of the scene. We render the reconstruction from this novel trajectory (in dotted-orange), and use the rendering to identify co-visible pixels in the original source view (in orange). The source video and its masked variant are used to produce a self-supervised training pair for training CogNVS, our "3D-aware" video inpainting diffusion model.
We propose to train CogNVS in a self-supervised manner. This allows us to use a large corpus of 2D videos. For each casually captured monocular video $\mathbf{V}_{\rm src}$, we obtain its 3D reconstruction $\mathcal{G}_{\rm src}$ and odometry $\mathbf{c}_{\rm src}$ from off-the-shelf SLAM frameworks [10]. As demonstrated in Fig. 3, we sample $N$ arbitrary camera trajectories in order to create training pairs from 2D videos, described as follows.
We first obtain the “center” of the scene by considering the pixel at the optical center in the first frame of the given video, similar to a prior work [51]. We then construct a bounded region $\mathcal{S}$ in spherical coordinates, around the camera center of $c_{\rm src}^{1}$. Within this region, we uniformly sample start and end spherical coordinates of each new camera trajectory, and then again sample two intermediate camera locations between the start and end spherical coordinates to ensure smoothness during interpolation. Camera poses are obtained by converting the spherical coordinates into euclidean space to get translations, and camera rotations are obtained such that the look-at vector always points to the center of the scene. Using the four sampled camera poses, we do bicubic interpolation on the $\rm SE(3)$ manifold. This results in a set of smooth camera trajectories, $\{\mathbf{c}_{\rm n}\}_{n=1}^{N}$ which are then used to construct the training pairs. With $N$ trajectories, we can obtain “partial” novel-view renderings as,
@@
[…]

### v2

# Reconstruct, Inpaint, Test-Time Finetune: Dynamic Novel-view Synthesis from Monocular Videos

@@ Abstract
Figure 1: We present CogNVS, a video diffusion model that enables novel-view synthesis of dynamic scenes. Given an in-the-wild monocular video of a dynamic scene, we first reconstruct the scene, render it from the target novel-view and inpaint any unobserved regions. Because CogNVS can be pre-trained via self-supervision, it can also be test-time-finetuned on a given target video, enabling it to zero-shot generalize to novel domains. Our simple pipeline outperforms almost all prior state-of-the-art for dynamic novel-view synthesis. We show outputs from CogNVS from two unseen videos; a real-world video above, and a generated video below.
@@ 1 Introduction
Rapid advances in static 3D scene representations Mildenhall et al. (2020); Kerbl et al. (2023) have paved the way for spacetime understanding of the dynamic world. This has enabled photorealistic content creation and immersive virtual reality applications. In this work, we focus on the problem of novel-view synthesis from casually-captured monocular videos of dynamic scenes.
@@
Prior work on dynamic view synthesis addresses this task from two extremes. The first class of methods “test-time” optimize a new 4D representation from scratch for every new test video. While this ensures physically-plausible scene geometry, careful choices in modeling scene motion – in the form of an independent deformation field, or learnable temporal offsets – have to be made Wang et al. (2024); Li et al. (2023); Wu et al. (2024a). More importantly, it can take on the order of hours to optimize and render a novel-view video. An attractive alternative is to train large feed-forward video models directly for view synthesis Zhang et al. (2024b); Hong et al. (2024). While inference on such models is dramatically faster (on the order of milliseconds), the resulting renderings often are not as accurate as their test-time optimized counterparts. From a pragmatic perspective, such models need to be trained on mega-scale multi-view training data, which is difficult to obtain for dynamic scenes.
@@
addresses the above challenges by decomposing the problem of dynamic view-synthesis into three distinct stages. First, we lean on the success of non-rigid structure from motion Lu et al. (2024); Zhang et al. (2024a); Li et al. (2024b) approaches that produce reconstructions of visible scene regions, sometimes known as "2.5D" reconstructions (since occluded regions are not reconstructed). We point out that such reconstructions can be trivially produced for casual mobile videos captured with depth sensors and egomotion Gao et al. (2022). When such reconstructions are rendered from a target novel view, previously-hidden regions will not not be rendered. To “inpaint” these regions, we train a 2D video-inpainter – CogNVS – by fine-tuning a video diffusion model (CogVideoX Yang et al. (2024c)) to condition on the partially-observable novel-view pixels. Importantly, we allow CogNVS to also update the appearance of previously-visible pixels, allowing our pipeline to model view-dependent (dynamic) scene effects.
The key insight of our work is that CogNVS can be trained on any 2D video via self-supervision. However, rather than training our inpainter with random 2D masks, we make use of 3D multi-view supervision that better captures 3D scene visiblity, similar to prior art Weber et al. (2024). Specifically, given a 2D training video, we first reconstruct it (with an off-the-shelf method such as MegaSAM) and then render the reconstruction from a random camera trajectory. This rendering is used to identify co-visible pixels from the source video that remain visible in the novel views. This original source video and its co-visible-only masked variant can now form a training pair for 3D-consistent video inpainting. Importantly, because such a training pair does not require ground-truth 3D supervision, CogNVS can be trained on diverse in-the-wild 2D videos. We use dynamic scenes from TAO Dave et al. (2020), SA-V Ravi et al. (2024), Youtube-VOS Xu et al. (2018), and DAVIS Pont-Tuset et al. (2017). Equally as important, we use the same paradigm to test-time finetune CogNVS on the test video-of-interest. We show that this allows our pipeline to “zero-shot” generalize to test videos that were never seen during training. We argue that our test-time finetuning of 2D diffusion models can be seen as the “best-of-both-worlds”, by leveraging large-scale training data (for data-driven robustness) and test-time optimization (for accuracy).
@@ 2 Related Work
has seen recent advancements with the rise of implicit scene representations like NeRFs Mildenhall et al. (2020) and Gaussian primitives Kerbl et al. (2023); Keselman and Hebert (2022); Chen et al. (2024). We have seen widespread efforts in scaling these representations to model larger scenes Turki et al. (2022); Xiangli et al. (2022); Tancik et al. (2022), making them faster to fit Lee et al. (2025); Garbin et al. (2021); Fan et al. (2024); Kerbl et al. (2023); Müller et al. (2022); Cao and Johnson (2023); Deng et al. (2022); Chen et al. (2021), anti-aliased Liang et al. (2024); Barron et al. (2021, 2022); Hu et al. (2023); Barron et al. (2023), and extend to representing dynamic scenes Stearns et al. (2024); Gao et al. (2024); Pumarola et al. (2021); Lee et al. (2025). The most popular paradigms have been the adoption of dynamic NeRFs Mildenhall et al. (2020); Pumarola et al. (2021); Gao et al. (2021); Park et al. (2021) and deformable Gaussian primitives Kerbl et al. (2023); Wu et al. (2024a); Yang et al. (2024b); Luiten et al. (2024) for modeling scene dynamics, apart from using voxel grids Khurana et al. (2023, 2022) or learnable tokenzation Zhang et al. (2023). Most approaches need multi-view posed videos as input, and only recently monocular view synthesis has gained traction Gafni et al. (2021); Lei et al. (2024); Wang et al. (2024); Li et al. (2021). However, each of the aforementioned approaches have to be test-time optimized separately for every new test video, are slow to optimize and yet fail to recover highly-detailed dynamic scene content Lei et al. (2024). Moreover, there is no focus on predicting the unobservable scene content, which is exacerbated by benchmarking metrics that only evaluate co-visible pixels Gao et al. (2022) in training and inference views and therefore encourage benchmarking on novel views that are not too far apart from the training views. Our approach instead reformulates dynamic view synthesis as an inpainting task, which specifically focuses on generating parts of the scene that were occluded from the training views, thereby facilitating extreme novel view synthesis for dynamic scenes. Our large-scale pretraining for feed-forward novel-view inpainting enables data-driven robustness.
@@
Figure 2: CogNVS overview. During training (left), given a 2D source video (in blue) of a dynamic scene, we first reconstruct the scene using off-the-shelf monocular reconstruction algorithms like MegaSAM Li et al. (2024b) to obtain the 3D scene geometry, $\mathcal{G}_{\rm src}$ and camera odometry, $\mathbf{c}_{\rm src}$. We then sample a set of arbitrary camera trajectories $\{\mathbf{c}_{1},\cdots,\mathbf{c}_{N}\}$ to simulate plausible occluded geometries, $\{\mathcal{G}^{\rm cov}_{{\rm src},1},\cdots,\mathcal{G}^{\rm cov}_{{\rm src},N}\}$ which when rendered from original camera trajectory, $\mathbf{c}_{\rm src}$ produces a mask of source pixels that are co-visible in the sampled trajectory (in orange). The source video and its masked variant produce a self-supervised training pair for learning CogNVS, our video inpainting diffusion model (visualized in Fig. 3). At inference (right), we finetune CogNVS on the given input sequence by similarly constructing self-supervised training pairs. The final novel-view is then generated using the finetuned CogNVS in a feed-forward manner.
@@
approaches have emerged Van Hoorick et al. (2024); Wu et al. (2024b); Ren et al. (2025); YU et al. (2025); Yu et al. (2024b); Liu et al. (2023b); Yang et al. (2024a) which train for view synthesis in a feed-forward manner with large-scale data. One class of methods is based on transformer architectures, more often than not trained with multi-view supervision and rendering in the loop Hong et al. (2024); Zhang et al. (2024b); Jiang et al. (2024); Wei et al. (2024); Xie et al. (2024); Ren et al. (2024). Another class of methods reformulate novel-view synthesis as a conditional generation task and use diffusion-based generative architectures Liu et al. (2023b); Shi et al. (2023) for the same. Initially, the focus was on developing data-driven pipelines for static novel-view synthesis Liu et al. (2023b, a, 2024); Yu et al. (2024b); Zhang et al. (2024b); Wei et al. (2024); Chan et al. (2023); Watson et al. (2023), or exploiting data-driven priors Poole et al. (2023); Wang et al. (2023); Sun et al. (2024); Chen et al. (2025); Tang et al. (2023); Khurana and Ramanan (2024); Raj et al. (2023); You et al. (2025), using multi-view posed image inputs. However, the focus is now shifting to dynamic view synthesis of casually captured videos in an unconstrained setting Van Hoorick et al. (2024); Ren et al. (2025); YU et al. (2025); Yang et al. (2024a); Li et al. (2024a); Jeong et al. (2024). These approaches allow a greater level of hallucination of unseen scene components which instills the capability of view synthesis for camera poses that are far apart from the training views. We fit into this setting. While the data-driven learning provides faster inference times and a broad generalization, it compromises on 3D geometric accuracy and physical-plausibility which introduces unrealistic artifacts in the synthesized outputs (e.g., objects suddenly exist or cease to exist) YU et al. (2025). In this work, we highlight that test-time finetuning is crucial to preserving the 3D geometry of the scene and reducing implausible artifacts.
@@
is a long-standing paradigm to curb distribution shifts in machine learning algorithms and improve their generalization. It’s origin lies in early-age algorithms for optical character recognition Bottou and Vapnik (1992) and text classification Joachims and others (1999), where the algorithm adjusts itself after observing the test data. A decade back, it popularly resurfaced for super-resolution Shocher et al. (2018) where learning to super-resolve an image was achieved by downsampling and super-resolving the test image. Domain generalization approaches for vision Sun et al. (2020); Chen et al. (2022); Iwasawa and Matsuo (2021); Chen et al. (2023); Gandelsman et al. (2022); Zhao et al. (2024) soon took inspiration from this breakthrough and recently, chain-of-thought prompting Wei et al. (2022) and general LLM reasoning Akyürek et al. (2024) in natural language processing adapted this paradigm. The most recent adoption was seen in 4D reconstruction and tracking Feng et al. (2025), and we similarly explore this paradigm further in our work.
@@ 3 Method
Given a monocular video of a dynamic scene, $\mathbf{V}_{\rm src}=\{\mathbf{V}_{\rm src}^{t}\}_{t=1}^{T}$, we want to generate a novel view of the observed scene, $\mathbf{V}_{\rm nvs}=\{\mathbf{V}_{\rm nvs}^{t}\}_{t=1}^{T}$ from a target camera pose. As discussed (c.f. Fig. 2), we achieve this by decomposing the task into three distinct stages – (1) obtain an off-the-shelf reconstruction of the observed scene over time, (2) render the scene from the novel views and inpaint the non-co-visible regions, and (3) curb the train-test distribution shift with test-time finetuning. Note that the first two stages of our pipeline are similar to concurrent work YU et al. (2025); Ren et al. (2025) but importantly we find that test-time finetuning is a crucial stage to allow generalization. We now describe each of the stages in detail.
@@ 3.1 Dynamic view synthesis as structured inpainting
We use off-the-shelf SLAM frameworks, like MegaSAM Li et al. (2024b), to obtain a reconstruction of the given scene. Formally, let the underlying 3D structure of the world as observed by $\mathbf{V}_{\rm src}$ be represented by, $\mathcal{G}_{\rm src}=\{\mathbf{X}_{\rm src}^{t}\}_{t=1}^{T}$, where $\mathbf{X}_{\rm src}^{t}$ are the evolving 3D primitives (points, Gaussians, etc.) across time, $t$. Any physical properties of the primitives are omitted from this discussion for simplicity. Let the recovered camera poses from which $\mathbf{V}_{\rm src}$ was observed be, $\mathbf{c}_{\rm src}=\{\mathbf{c}_{\rm src}^{t}\}_{t=1}^{T}$, where $\mathbf{c}$ denotes a camera pose and is formulated as, $\mathbf{c}=(\mathbf{R},\mathbf{t})\in\mathrm{SE}(3)$ lie group. The source video $\mathbf{V}_{\rm src}$ can be obtained by using a rendering function $\mathcal{R}$ as,
@@
For obtaining $\mathbf{V}_{\rm nvs}$, we note that a subset of 3D primitives that must be visible from $\mathbf{c}_{\rm src}$, are already available in the reconstructed scene geometry, $\mathcal{G}_{\rm src}$. Therefore, a partial observation of the world in the form of co-visible pixels Gao et al. (2022) from novel views, $\mathbf{c}_{\rm nvs}=\{c_{\rm nvs}^{t}\}_{t=1}^{T}$, can be rendered as follows,
@@
At this point, the novel view synthesis is incomplete, and all missing regions have to be generated. To this end, we train a conditional video diffusion model, CogNVS (denoted by $\epsilon_{\theta}$) built on top of a recently proposed transformer-based video diffusion model Yang et al. (2024c). CogNVS takes in the partially observed novel view video and generates an inpainted novel-view of the scene. The overall CogNVS pipeline first employs a 3D causal VAE to compress the conditioning $\mathbf{V}_{\rm nvs}^{\rm cov}$ and target novel-view $\mathbf{V}_{\rm src}$ into latent representations $\mathbf{z}_{\rm cond}=\mathcal{E}(\mathbf{V}_{\rm nvs}^{\rm cov})$ and $\mathbf{z}_{0}=\mathcal{E}(\mathbf{V}_{\rm src})$ respectively, enabling efficient training while preserving temporal coherence and photometric fidelity. Here, $\mathcal{E}$ is the VAE encoder. Gaussian noise is then added to the target latent $\mathbf{z}_{0}$, and the resulting noisy latent is concatenated with the conditional latent $\mathbf{z}_{\rm cond}$. This joint representation is passed through a self-attention transformer equipped with 3D rotary positional embeddings (3D-RoPE) Su et al. (2024) and adaptive layer normalization, which predicts the added noise. The training objective follows a score matching formulation:
@@ 3.2 Data generation for self-supervised training
Figure 3: Self-supervised training data generation. To curate a large training set for video inpainting, we first reconstruct an input source 2D video (in blue) with an off-the-shelf monocular SLAM system. After reconstruction, we randomly sample NN pairs of ‘start’ and ‘end’ camera poses around a spherical region, $\mathcal{S}$ of the estimated camera pose in the given 2D video. $\mathcal{S}$ is bounded by a predefined deviation in the spherical coordinate axes, similar to a prior work Yu et al. (2024b). We sample a $\rm SE(3)$} camera trajectory that interpolates the start and end poses while looking at the center of the scene. We render the reconstruction from this novel trajectory (in dotted-orange), and use the rendering to identify co-visible pixels in the original source view (in orange). The source video and its masked variant are used to produce a self-supervised training pair for training CogNVS, our "3D-aware" video inpainting diffusion model.
We propose to train CogNVS in a self-supervised manner. This allows us to use a large corpus of 2D videos. For each casually captured monocular video $\mathbf{V}_{\rm src}$, similar to prior work Han et al. (2022); Lv et al. (2025); YU et al. (2025), we obtain its 3D reconstruction $\mathcal{G}_{\rm src}$ and odometry $\mathbf{c}_{\rm src}$ from off-the-shelf SLAM frameworks Li et al. (2024b). As demonstrated in Fig. 3, we sample $N$ arbitrary camera trajectories in order to create training pairs from 2D videos, described as follows.
We first obtain the “center” of the scene by considering the pixel at the optical center in the first frame of the given video, similar to a prior work YU et al. (2025). We then construct a bounded region $\mathcal{S}$ in spherical coordinates, around the camera center of $c_{\rm src}^{1}$. Within this region, we uniformly sample start and end spherical coordinates of each new camera trajectory, and then again sample two intermediate camera locations between the start and end spherical coordinates to ensure smoothness during interpolation. Camera poses are obtained by converting the spherical coordinates into euclidean space to get translations, and camera rotations are obtained such that the look-at vector always points to the center of the scene. Using the four sampled camera poses, we do bicubic interpolation on the $\rm SE(3)$ manifold. This results in a set of smooth camera trajectories, $\{\mathbf{c}_{\rm n}\}_{n=1}^{N}$ which are then used to construct the training pairs. With $N$ trajectories, we can obtain “partial” novel-view renderings as,
@@
$\mathbf{V}_{{\rm src},n}^{\rm cov}=\mathcal{R}\bigl(\mathcal{G}^{\rm cov}_{{\rm src},n},\mathbf{c}_{\rm src}\bigr)\hskip 19.91684pt{\rm s.t.}\hskip 19.91684pt\mathcal{D}=\{(\mathbf{V}_{{\rm src},n}^{\rm cov},\mathbf{V}_{\rm src})\}\forall n\in[1,N]$